# thanks: These two authors contribute equally in this workthanks: These two authors contribute equally in this work

### v1

@@ Abstract
Topological superconductivity.—We first focus on the SC phase in zero or small displacement field regime. A comparison of the energy gain $\Delta E$ between possible SC orders and $120^{\circ}$ AFM order is shown in Fig. 4(a), which indicates that the system is in SC phase for $\mathcal{V}_{z}<\mathcal{V}_{z,c}\approx 4.0$ meV ($5.2$ meV) for $V=10$ meV ($12.5$ meV), qualitatively consistent with the critical field $\mathcal{V}_{z,c}^{\text{exp}}\approx 2.6$ meV observed in the experiment [24]. The pairing of this SC phase is chiral with mixed $p_{x}\pm ip_{y}$ and $d_{xy}\mp id_{x^{2}-y^{2}}$-wave in the $E$ representation of $C_{3v}$, spontaneously breaking the time reversal symmetry $\mathcal{T}$. More importantly, such chiral SC phase, which fits into the A-class (namely, breaking $\mathcal{T}$ but preserving $S_{z}$) of the Altland-Zirnbauer tenfold classification scheme [52, 53], is topologically non-trivial with Chern number computed to be $C=\pm 2$, suggesting tWSe2 as a promising candidate for realizing chiral topological SC. The internal structure of the SC order as a function of $\mathcal{V}_{z}$ for $V=10$ meV is also illustrated in Fig. 4(b), where the consistency constraint between time reversal symmetry $\mathcal{T}$ and $C_{3v}$ mirror plane symmetry enables us to fix a simple gauge such that all singlet pairings $\Delta^{S}_{\alpha}(\delta)$ are real while all triplet pairings $\Delta^{T}_{\alpha}(\delta)$ are imaginary. It is worth emphasizing that, unlike the moiré Hubbard model [42, 43, 16, 17, 18], the mixing of singlet and triplet pairings are allowed for all $\mathcal{V}_{z}$, especially for $\mathcal{V}_{z}=0$ due to the absence of emergent spin-valley SU(2) symmetry [54], and the dominance of $d_{xy}\mp id_{x^{2}-y^{2}}$ component over the $p_{x}\pm ip_{y}$ one in the SC phase is in accordance with the result of Chern number ±$\pm 2$. Moreover, the SC is enhanced (suppressed) on the B (A) sublattice upon increasing the displacement field $\mathcal{V}_{z}$ from zero, closely following the trend of Fermi surface DOS in Fig. 2(b).
@@
Antiferromagnetic correlated insulators.—As the displacement field is further increased beyond a critical field $\mathcal{V}_{z,c}$, the system transitions into the $120^{\circ}$ AFM phase. To further determine its transport properties, we compute the AFM mean-field charge gap for $\tilde{\nu}=-3$. As shown in Fig. 4(c), a broad intermediate range of AFM insulator (AFM-I) phase appears for $\mathcal{V}_{z,c}\lesssim\mathcal{V}_{z}<\mathcal{V}^{\prime}_{z,c}\approx 20.8$ meV and an AFM metal (AFM-M) phase for $\mathcal{V}_{z}>\mathcal{V}^{\prime}_{z,c}$, closely matching the phenomenology of the experiment [24]. In the mean-field framework, depending on the value of $V$ employed in the model, the SC to AFM-I transition either exhibits a tiny intermediate AFM-M phase as shown in Fig. 4(e) or occurs as a direct first-order transition as shown in Fig. 4(f). To explain the transport evidences of continuous superconductor-insulator transition [24], disorder could play an important role. The continuous transition into the SC phase might potentially be a percolation transition [55, 56] of local SC regions induced by disorders or a disordered-round first-order transition. We also present the magnitude of AFM orders on different sublattices as a function of $\mathcal{V}_{z}$ in Fig. 4(d). Except for small $\mathcal{V}_{z}$, where the AFM order is stronger on B sublattice due to its larger DOS (see Fig. 2(b)), the AFM order generally favors A sublattice, as the holes are mostly concentrated there in response to the displacement field $\mathcal{V}_{z}$, as well as its better approximation for commensurate nesting wave vector $\bm{Q}$. And the sudden drop of AFM orders at large $\mathcal{V}_{z}$ regime coincides with the disappearance of B sublattice hole pockets at $\kappa_{\pm}$ points and the sharp decline in B sublattice DOS, as illustrated in Fig. 2, indicating that the holes on B sublattice play a crucial role in mediating the AFM order.
@@
Summary and Discussions.— In summary, we have constructed a three-band tight-binding model through direct Wannierization, and incorporated onsite Hubbard repulsion and NNN attraction to explain the SC and correlated insulator phase observed in the $3.65^{\circ}$ tWSe2 at filling $\nu=-1$ [24]. Our mean-field analysis indicates that, the SC phase is an A-class topological SC with Chern number $C=\pm 2$, featuring the inter-valley mixed $p_{x}\pm ip_{y}$ and $d_{xy}\mp id_{x^{2}-y^{2}}$-wave pairing symmetry, and the correlated insulator phase is explained by the $120^{\circ}$ AFM order. The topological band structure in our model can be crucial for understanding how SC arises for flat band systems, since non-trivial lower bounds of SC superfluid weight exists [57, 58, 59] due to quantum geometric effects, which deserves more detailed future theoretical studies. Moreover, our results also suggest that the tWSe2 homobilayers could provide new possibilities for realizing topological SC, which call for more detailed experimental studies to further uncover its nature. Identifying the topological edge states by measuring quantized thermal Hall conductance [60], or verifying the chiral nature of the SC pairing symmetry through phase sensitive Josephson junction [61], will surely open new opportunities in twisted TMD systems.
@@
Acknowledgements—We would like to thank Andrei Bernevig for helpful discussions. This work is supported in part by the Innovation Program for Quantum Science and Technology (grant No. 2021ZD0302502), by MOSTC under Grant No. 2021YFA1400100, by NSFC under Grant Nos. 12347107 and 12334003. HY acknowledgement the support in part by the Xplorer Prize through the New Cornerstone Science Foundation.
@@ I Hopping Parameters of the Tight-Binding Model
In this section, we provide additional information of the tight-binding model we adopted in the main text, derived from direct Wannierization of the continuum model for $3.65^{\circ}$ tWSe2 at $V_{z}=0$. Fig. S1 illustrates the symmetry inequivalent representative hopping bonds up to $5$th-nearest-neighbor, and we provide the hopping parameters for $K$ valley on these representative bonds as well as the onsite potentials as:
$t_{0}^{A}=t_{0}^{B}=-15.26\text{ meV},\quad t_{0}^{C}=-42.89\text{ meV}$ (S1)
$t^{BC}_{1}=7.12\text{ meV},\quad t^{CC}_{\sqrt{3}}=2.76\text{ meV},\quad t^{AC}_{2}=1.15\text{ meV}$
$t^{AC}_{\sqrt{7}}=3.01e^{i0.965\pi}\text{ meV},\quad t^{CC}_{3}=0.43\text{ meV}$
$t^{AB}_{1}=-4.99\text{ meV},\quad t^{BB}_{\sqrt{3}}=5.16e^{i0.618\pi}\text{ meV}$
$t^{AB}_{2}=0.7\text{ meV},\quad t^{AB}_{\sqrt{7}}=0.78\text{ meV},\quad t^{BB}_{3}=-0.31\text{ meV}$
@@
There are several remarks on the symmetry properties of these hopping parameters. First, hopping parameters related by $C_{3z}$ symmetry will (will not) be identical if $\alpha,\beta$ have identical (different) $C_{3z}$ eigenvalues (see main text). Second, $C_{2y}\mathcal{T}$ and $\mathcal{I}$ symmetries constrain most but not all hopping parameters to be real. Third, the hopping parameters for $K$ and $-K$ valley are related by time reversal symmetry $\mathcal{T}$. Last, unlike the single-band moiré Hubbard model, where emergent spin-valley SU(2) symmetry exists at $V_{z}=0$, the spin-valley symmetry here remains U(1) due to the presence of complex hopping parameters.
Figure S1: Illustration of the representative hopping bonds up to $5$th-nearest-neighbor.
II Derivations and Supplemental Results of Superconducting Mean-Field Analysis

### v2

@@ III Results
Topological superconductivity.—We first focus on the SC phase in the zero- or small-displacement field regime. A comparison of the energy gain $\Delta E$ between possible SC orders and the $120^{\circ}$ AFM order is shown in Fig. 4(a), which indicates that the system is in the SC phase for $\mathcal{V}_{z}<\mathcal{V}_{z,c}\approx 2.2$ meV ($3.4$ meV) for $V_{2}=10$ meV ($12.5$ meV), qualitatively consistent with the critical field $\mathcal{V}_{z,c}^{\text{exp}}\approx 2.6$ meV observed in the experiment [71]. The pairing of this SC phase is chiral with mixed $d_{xy}\pm id_{x^{2}-y^{2}}$ and $p_{x}\mp ip_{y}$ wave in the $E$ representation of $C_{3v}$, spontaneously breaking the time-reversal symmetry $\mathcal{T}$ [105]. More importantly, such a chiral SC phase, which fits into the A-class (namely, breaking $\mathcal{T}$ but preserving the $S_{z}$) of the Altland-Zirnbauer tenfold classification scheme [97, 98, 99], is topologically non-trivial with Chern number computed to be $C=\pm 2$, suggesting tWSe2 as a promising candidate for realizing chiral topological SC. The internal structure of the SC order as a function of $\mathcal{V}_{z}$ for $V_{2}=10$ meV is also illustrated in Fig. 4(b), where the consistency constraint between time-reversal symmetry $\mathcal{T}$ and $C_{3v}$ mirror plane symmetry enables us to fix a simple gauge such that all singlet pairings $\Delta^{S}_{\alpha}(\delta)$ are real while all triplet pairings $\Delta^{T}_{\alpha}(\delta)$ are imaginary. It is worth emphasizing that, unlike the single-band moiré Hubbard model [49, 73, 74, 75, 60, 61, 67, 63, 64, 62, 65, 66, 76, 100, 101], the mixing of singlet and triplet pairings is allowed for all $\mathcal{V}_{z}$ in our three-band model, especially for $\mathcal{V}_{z}=0$, where the emergent spin-valley SU(2) symmetry is absent and the pseudo-inversion symmetry $\mathcal{I}$ does not forbid such mixing. Also, the dominance of the $d_{xy}\pm id_{x^{2}-y^{2}}$ component over the $p_{x}\mp ip_{y}$ one in the SC phase is in accordance with the result of Chern number $\pm 2$. Moreover, SC is enhanced (suppressed) on the B (A) sublattice upon increasing the displacement field $\mathcal{V}_{z}$ from zero, closely following the trend of the Fermi surface DOS in Fig. 2(b).
@@
Antiferromagnetic correlated insulators.—As the displacement field is further increased beyond a critical field $\mathcal{V}_{z,c}$, the system transitions into the $120^{\circ}$ AFM phase. To further determine its transport properties, we compute the AFM mean-field charge gap for $\tilde{\nu}=-3$. As shown in Fig. 4(c), a broad intermediate range of the AFM insulator (AFM-I) phase appears for $\mathcal{V}_{z,c}\lesssim\mathcal{V}_{z}<\mathcal{V}^{\prime}_{z,c}\approx 21.5$ meV and an AFM metal (AFM-M) phase for $\mathcal{V}_{z}>\mathcal{V}^{\prime}_{z,c}$, closely matching the phenomenology of the experiment [71]. In the mean-field framework, depending on the value of $V_{2}$ employed in the model, the SC to AFM-I transition either exhibits a tiny intermediate AFM-M phase as shown in Fig. 4(e) or occurs as a direct first-order transition as shown in Fig. 4(f). To explain the transport evidence of continuous superconductor-insulator transition [71], disorder could play an important role. The continuous transition into the SC phase might potentially be a percolation transition [106, 107] of local SC regions induced by disorders or a disorder-rounded first-order transition. We also present the magnitude of AFM orders on different sublattices as a function of $\mathcal{V}_{z}$ in Fig. 4(d). Except for small $\mathcal{V}_{z}$, where the AFM order is stronger on the B sublattice due to its larger DOS (see Fig. 2(b)), the AFM order generally favors the A sublattice, as the holes are mostly concentrated there in response to the displacement field $\mathcal{V}_{z}$, as well as its better approximation for the commensurate nesting wave vector $\bm{Q}$. And the sudden drop of AFM orders at large $\mathcal{V}_{z}$ regime coincides with the disappearance of B sublattice hole pockets at $\kappa_{\pm}$ points and the sharp decline in B sublattice DOS, as illustrated in Fig. 2, indicating that the holes on B sublattice play a crucial role in mediating the AFM order.
IV Discussion
In summary, we have constructed a three-band tight-binding model through direct Wannierization, and incorporated onsite Hubbard repulsion and NNN attraction to explain the SC and correlated insulator phase observed in the $3.65^{\circ}$ tWSe2 at filling $\nu=-1$ [71]. Our mean-field analysis indicates that, the SC phase is an A-class topological SC with Chern number $C=\pm 2$, featuring the inter-valley mixed $d_{xy}\pm id_{x^{2}-y^{2}}$ and $p_{x}\mp ip_{y}$-wave pairing symmetry, and the correlated insulator phase is explained by the $120^{\circ}$ AFM order. We further demonstrate that the metallic behavior when the filling is away from $\nu=-1$ can also be qualitatively understood within our mean-field framework (see Sec. VIII of the Supplementary Information [104]). Compared with the $5^{\circ}$ tWSe2 system [72], the experimental features show notable similarities despite differences in low-energy band structure and interaction strength [83]. Recent theoretical advances [108] suggest the possibility of a unifying underlying mechanism, motivating future efforts toward a unified description.
@@ IV Discussion
The topological band structure in our model can be crucial for understanding how SC arises for flat band systems, since non-trivial lower bounds of SC superfluid weight exist [109, 110, 111, 112, 113, 114] due to quantum geometric effects, which deserves more detailed future theoretical studies. Moreover, our results also suggest that the tWSe2 homobilayers could provide new possibilities for realizing topological SC, which call for more detailed experimental studies to further uncover its nature. Identifying the topological edge states by measuring quantized thermal Hall conductance [115, 116, 117], or verifying the chiral nature of the SC pairing symmetry through the phase-sensitive Josephson junction [118, 119, 120], will surely open new opportunities in twisted TMD systems.
@@
V Methods
We construct a three-band tight-binding model by a direct Wannierization of the continuum model of $3.65^{\circ}$ twisted bilayer WSe2, with the resulting hopping parameters given in Supplementary Information Sec. I. The model incorporates onsite Hubbard repulsion and nearest- or next-nearest-neighbor attractions. A detailed estimation of the Hubbard interaction parameters is provided in Supplementary Information Sec. III and Sec. IV, while a microscopic explanation of the attractive terms is presented by the random phase approximation calculation in Supplementary Information Sec. V. The interacting model is analyzed within the standard mean-field framework, with full derivations and self-consistent procedures provided in Supplementary Information Sec. VI and Sec. VII.
VI Data Availability
Source data for all figures in the main article are available in the Supplementary Data file.
VII Code Availability
The codes used in this study are available from the corresponding author upon request.
@@ I Hopping Parameters of the Tight-Binding Model
In this section, we provide additional information of the tight-binding model hopping parameters we adopted in the main text, derived from direct Wannierization of the continuum model for $3.65^{\circ}$ tWSe2 at $V_{z}=0$. Fig. S1 illustrates the symmetry inequivalent representative hopping bonds up to $5$th-nearest-neighbor, and we provide the hopping parameters for $K$ valley on these representative bonds as well as the onsite potentials as:
[…]
@@
There are several remarks on the symmetry properties of these hopping parameters. First, hopping parameters related by $C_{3z}$ symmetry will (will not) be identical if $\alpha,\beta$ have identical (different) $C_{3z}$ eigenvalues (see main text). Second, $C_{2y}\mathcal{T}$ and $\mathcal{I}$ symmetries constrain most but not all hopping parameters to be real. Third, the hopping parameters for $K$ and K-K valley are related by time-reversal symmetry $\mathcal{T}$. Last, unlike the single-band moiré Hubbard model, where emergent spin-valley SU(2) symmetry exists at $V_{z}=0$, the spin-valley symmetry here remains U(1) due to the presence of complex hopping parameters.
Figure S1: Illustration of the representative hopping bonds up to 55th-nearest-neighbor.
II Displacement Field Dependence of the Wannier Functions
The layer-polarization Wannierization scheme can be applied not only to tWSe2 at zero displacement field, but also at finite displacement fields. Fig. S2 illustrates the density distribution of Wannier functions obtained from the continuum model with the displacement field set to $V_{z}=20$ meV (see Eq. (2) in the main text).
Compared with the Wannier functions at zero displacement field (Fig. 1(b) in the main text), we find that the A and B Wannier functions remain almost unchanged (A (B) becomes slightly more localized (extended)) and remain strongly polarized in opposite layers. In contrast, the C Wannier function exhibits a more noticeable redistribution though remaining layer hybridized. This behavior can be simply understood as follows: the A and B Wannier functions are approximate eigenstates of the displacement field due to their strong layer polarization, while the C Wannier function is not. Nevertheless, we neglect this modification of the C orbital in the main text, as it lies far from the Fermi surface and does not play a significant role in the low-energy physics.
To further justify this approximation, we compute the onsite chemical potential terms $t_{0}^{\alpha}$ of the Wannier functions at $V_{z}=20$ meV: $t_{0}^{A}=-7.77$ meV, $t_{0}^{B}=-22.72$ meV, and $t_{0}^{C}=-41.10$ meV. These results confirm that the C orbital remains energetically well separated from the low-energy physics. Moreover, we can extract the effective displacement field $\mathcal{V}_{z}=t_{0}^{A}-t_{0}^{B}=14.95$ meV in our tight-binding model (see Eq. (5) in the main text). The fact that $\mathcal{V}_{z}$ is slightly smaller than the original $V_{z}$ can be attributed to the A and B Wannier functions not being perfectly layer polarized. These analyses confirm the robustness of the assumptions underlying our model.
Figure S2: Density distribution of Wannier functions on two layers at $V_{z}=20$ meV, with unit cell shown as dashed line.
III Mean-field Results with Different Interacting Parameters
In the main text, we adopt the interacting parameters $U_{A}=U_{B}=U_{C}=37.5$ meV, $V_{2}=10$ meV or $12.5$ meV as representative parameters of $3.65^{\circ}$ tWSe2. This section further explores our mean-field model over a broad range of interacting parameters, offering valuable physical insights into the tWSe2 system and providing justifications of the parameter choices.
We start with the discussion of onsite Hubbard interaction. We emphasize that, although symmetries only impose $U_{A}=U_{B}$ while allowing $U_{C}$ to differ, the parameter choice can be simplified by setting $U_{A}=U_{B}=U_{C}=U$, as the C orbital is far from the Fermi surface (see Fig. 2 in the main text). Such simplification reduces independent tuning parameters without altering the essential physical properties, making the qualitative physics more transparent. Fig. S3 (a,b) illustrates the AFM energy gain per hole and the mean-field charge gap with $U_{A}=U_{B}=37.5$ meV with varying $U_{C}$, showing only minor differences (especially for the small displacement field regime close to the SC phase) even when $U_{C}$ is tuned over a relatively large range. To gain further insights into the strength of the Hubbard interaction, we then keep $U_{A}=U_{B}=U_{C}=U$ and vary the Hubbard $U$ from $25$ meV to $45$ meV. The results are shown in Fig. S3 (c,d), where both the AFM energy gain and the charge gap increase with $U$ as expected. Based on these results, we conclude that $U$ in the range of $35\sim 40$ meV is appropriate, where smaller $U$ weakens the insulating behavior while larger $U$ stabilizes AFM ground state over SC for small displacement field.
We now briefly comment on the values of the NNN attraction $V_{2}$. While an infinitesimal attraction $V_{2}$ should be sufficient to induce SC tendency (due to the Cooper instability), a relatively large $V_{2}=10$ or $12.5$ meV is adopted in the main text. This choice aims to provide a better explanation for the continuous SC-insulator transition in the experiment. In the main text, we attribute such continuous transition to disorder, which requires our mean-field results (without disorder) either exhibit a weak first-order direct SC-AFM insulator transition ($V_{2}=12.5$ meV), or feature a small intermediate AFM metal phase ($V_{2}=10$ meV). A weaker $V_{2}$ results in extended intermediate phase, while a larger $V_{2}$ will cause strong direct first-order transition, which is challenging to make continuous even when disorder is considered.
Figure S3: (a,b) The AFM energy gain per hole and the mean-field charge gap with $U_{A}=U_{B}=37.5$ meV and different $U_{C}$. (c,d) The AFM energy gain per hole and the mean-field charge gap with various $U_{A}=U_{B}=U_{C}=U$.
IV Interacting Parameters Directly from Wannier Functions
In addition to determining the interacting parameters by comparing the mean-field results with the experiment, these parameters can also be directly obtained from the Wannier functions. Given the considerable uncertainties in the microscopic mechanism of attractions, here we primarily focus on the more well-understood dual-gate screened Coulomb repulsion:
$\begin{split}H_{\text{dual-gate}}=\frac{1}{2A}\sum_{k,k^{\prime},q,l,l^{\prime},\tau,\tau^{\prime}}V_{0}(q)&\,c^{\dagger}_{k+ql\tau}c^{\dagger}_{k^{\prime}-ql^{\prime}\tau^{\prime}}c_{k^{\prime}l^{\prime}\tau^{\prime}}c_{kl\tau}\\ V_{0}(q)=\pi\xi^{2}V_{\xi}\frac{\tanh(\xi|q|/2)}{\xi|q|/2},&\quad V_{\xi}=\frac{e^{2}}{4\pi\epsilon_{0}\epsilon_{r}\xi}\end{split}$ (S2)
where $c_{kl\tau}$ annihilates a plane wave state with momentum $k$ in layer $l$ and spin/valley $\tau$, $A$ denotes the total area of the system, $\xi$ is the distance between the two metal gates, $e$ is the elementary charge, and $\epsilon_{0}$ ($\epsilon_{r}$) correspond to the vacuum (relative) dielectric constant.
Fig. S4 illustrates the interacting parameters obtained by expanding the dual-gate screened Coulomb interaction onto the Wannier functions, using an experimentally relevant gate distance of $\xi=10$ nm. Here, $V^{\bullet\alpha}_{d}$ denotes the interacting parameter between a Wannier function on $\alpha$ sublattice and another Wannier function at distance $d$ apart on the same representative bond shown in Fig. S1 (set NN with $d=1$, and $d=0$ corresponds to the onsite Hubbard interaction $U_{\alpha}$). The dielectric constant $\epsilon_{r}$ is determined by setting $U_{A/B}=37.5$ meV. We also plot the dual-gate screened Coulomb interaction in real space $V(r)$ (i.e., the Fourier transformation of $V_{0}(q)$), where the interacting parameters closely follow its trend at large distances. $V(r)$ can be written in summation form as:
$V(r)=V_{\xi}\sum_{n=-\infty}^{+\infty}\frac{(-1)^{n}}{\sqrt{(r/\xi)^{2}+n^{2}}}$ (S3)
which can be physically interpreted as the sum of all contributions from image charges induced by two metal gates.
Figure S4: Interacting parameters obtained by expanding the dual-gate screened Coulomb interaction with $\xi=10$ nm onto the Wannier functions. The black dashed line represents the dual-gate screened Coulomb interaction for two point charges.
Here, we offer some additional remarks on these interacting parameters. First, to ensure our mean-field results match the experiment, the dielectric constant is determined to be $\epsilon_{r}\sim 21.4$ (for $U_{A/B}=37.5$ meV, as in the main text). The relatively large value of $\epsilon_{r}$ can be attributed to additional screening effects in tWSe2 systems beyond dual-gate screening, including contributions from the dielectric environment or higher-energy bands. Second, we still adopt $U_{A}=U_{B}=U_{C}$ in the main text, though the Hubbard interaction $U_{C}$ is considerably smaller than $U_{A/B}$ in Fig. S4 (which can be naturally understood by the extension of the Wannier functions, see Fig. 1(b) in the main text). The main reason is that the Wannier functions obtained by the layer-polarization Wannierization scheme are known to be not maximally localized. In particular, the (not so localized) C Wannier function will become more localized if we further optimize its spread functional, resulting in a significantly larger $U_{C}$ (closer to $U_{A/B}$). Moreover, as already illustrated in Fig. S3(a,b), changing $U_{C}$ does not affect the main physics of the system, since the C Wannier function is away from the Fermi surface. Third, it is reasonable to retain only the onsite Hubbard interaction while neglecting further neighbor interactions. As shown in Fig. S4, interactions beyond NN are negligibly small, allowing us to safely discard them. Although the NN interaction (from directly expanding the dual-gate Coulomb interaction) has considerable strength (with $V^{AB}_{1}/U_{A/B}\sim 0.44$), we argue that the effective $V_{1}$ should be significantly reduced when accounting for possible attraction mechanisms or additional screening effects beyond dual-gate screening. Moreover, further minimization of the Wannier function spread can further enhance the Hubbard $U$. Therefore, our model retaining only onsite Hubbard repulsion should provide a good starting point for understanding the physical properties of tWSe2.
V Attractive Interactions from Random Phase Approximation Analysis
In this section, we provide a microscopic justification of the phenomenological attraction adopted in the main text, using a random phase approximation (RPA) analysis. The RPA method captures collective screening effects by summing the infinite series of bubble diagrams in the particle-hole channel, which can result in effective attractive interactions in certain real-space regions. In moiré superlattice systems, due to the small size of moiré Brillouin zone, Umklapp scattering must be treated explicitly in the RPA calculation, as we detailed below.
Here, we consider the RPA screening effects of dual-gate screened Coulomb interaction in Eq. (S2) projected to the topmost moiré band. The band projected density operator $\rho_{q+g}$ (where $q$ is in moiré Brillouin zone, $g$ is reciprocal wave vector) is given by:
$\rho_{q+g}=\sum_{k,\sigma}\Lambda^{k+q,k}_{g\sigma}c^{\dagger}_{k+q,\sigma}c_{k,\sigma}$ (S4)
where $\Lambda^{k+q,k}_{g\sigma}$ is the form factor of the projected density operator, defined by:
$\Lambda^{k_{1},k_{2}}_{g\sigma}=\sum_{g^{\prime},l}z^{*}_{g+g^{\prime},l,\sigma}(k_{1})z_{g^{\prime},l,\sigma}(k_{2})$ (S5)
[…]
There are several symmetry properties of the form factor $\Lambda^{k_{1},k_{2}}_{g\sigma}$ that can be useful:
$\Lambda^{k_{1},k_{2}}_{g\sigma}=\Lambda^{k_{1}+g^{\prime},k_{2}}_{g-g^{\prime}\sigma}=\Lambda^{k_{1},k_{2}+g^{\prime}}_{g+g^{\prime}\sigma}$ (S6)
$\Lambda^{k_{1},k_{2}}_{g\downarrow}=\left[\Lambda^{-k_{1},-k_{2}}_{-g\uparrow}\right]^{*}$ (S7)
$\Lambda^{k_{1},k_{2}}_{g\sigma}=\left[\Lambda^{k_{2},k_{1}}_{-g\sigma}\right]^{*}$ (S8)
where, Eq. (S6) requires the periodic gauge $z_{g}(k+g^{\prime})=z_{g+g^{\prime}}(k)$, and it is particularly useful for relating arbitrary momentum $k_{1},k_{2}$ back to the first Brillouin zone. Eq. (S7) requires the time-reversal symmetric gauge $z_{\downarrow}(k)=z_{\uparrow}^{*}(-k)$, which can relate the spin down quantities to the corresponding spin up ones.
Having projected the density operator to the topmost band, the particle-hole susceptibility $\Pi(i\omega_{m},q)_{g,g^{\prime}}$ is defined by the time-ordered density correlation function:
$\begin{split}\Pi(i\omega_{m},q)_{g,g^{\prime}}&=\frac{1}{A}\int_{0}^{\beta}d\tau e^{i\omega_{m}\tau}\int drdr^{\prime}e^{-i(q+g)r}e^{i(q+g^{\prime})r^{\prime}}\Pi(r,r^{\prime},\tau)\\ &=-\frac{1}{A}\int_{0}^{\beta}d\tau e^{i\omega_{m}\tau}\langle T_{\tau}\rho_{q+g}(\tau)\rho_{-q-g^{\prime}}(0)\rangle\end{split}$ (S9)
Since there are only discrete lattice translation symmetries $\Pi(r+a,r^{\prime}+a,\tau)=\Pi(r,r^{\prime},\tau)$ (where $a$ is some moiré lattice vector), momentum is only conserved modulo $g$. The above equation treats the Umklapp scattering structure explicitly and exactly.
Substitute the projected density operator Eq. (S4) into the above equation, we can evaluate the particle-hole susceptibility exactly:
$\begin{split}\Pi(i\omega_{m},q)_{g,g^{\prime}}&=-\frac{1}{A}\int_{0}^{\beta}d\tau e^{i\omega_{m}\tau}\sum_{kk^{\prime}\sigma\sigma^{\prime}}\Lambda^{k+q,k}_{g\sigma}\Lambda^{k^{\prime}-q,k^{\prime}}_{-g^{\prime}\sigma^{\prime}}\langle T_{\tau}c^{\dagger}_{k+q,\sigma}(\tau)c_{k,\sigma}(\tau)c^{\dagger}_{k^{\prime}-q,\sigma^{\prime}}(0)c_{k^{\prime},\sigma^{\prime}}(0)\rangle\\ &=\frac{1}{A}\int_{0}^{\beta}d\tau e^{i\omega_{m}\tau}\sum_{k\sigma}\Lambda^{k+q,k}_{g\sigma}\Lambda^{k,k+q}_{-g^{\prime}\sigma}G_{\sigma}(k,\tau)G_{\sigma}(k+q,-\tau)\\ &=\frac{1}{A}\sum_{k\sigma}\frac{n_{k\sigma}-n_{k+q\sigma}}{i\omega+\epsilon_{k\sigma}-\epsilon_{k+q\sigma}}\Lambda^{k+q,k}_{g\sigma}\Lambda^{k,k+q}_{-g^{\prime}\sigma}\\ &=\frac{1}{A}\sum_{k\sigma}\frac{n_{k\sigma}-n_{k+q\sigma}}{i\omega+\epsilon_{k\sigma}-\epsilon_{k+q\sigma}}\Lambda^{k+q,k}_{g\sigma}\left[\Lambda^{k+q,k}_{g^{\prime}\sigma}\right]^{*}\end{split}$ (S10)
where $G_{\sigma}(k,\tau)=-\langle T_{\tau}c_{\sigma}(\tau)c^{\dagger}_{\sigma}(0)\rangle$ is the single particle green’s function, $\epsilon_{k\sigma}$ is the band dispersion and $n_{k\sigma}$ is the Fermi-Dirac distribution.
Defining the bare interaction $V_{0}$ as a diagonal matrix of $g$:
$V_{0}(q)_{g,g^{\prime}}=V_{0}(q+g)\delta_{g,g^{\prime}}$ (S11)
Then, the RPA screened interaction follows directly from the standard Dyson series summation (summation over bubble diagrams), where all the quantities should be understood as matrix with index $g,g^{\prime}$:
$\begin{split}V^{RPA}(i\omega_{m},q)&=V_{0}(q)+V_{0}(q)\Pi(i\omega_{m},q)V_{0}(q)+V_{0}(q)\Pi(i\omega_{m},q)V_{0}(q)\Pi(i\omega_{m},q)V_{0}(q)+\cdots\\ &=[1-V_{0}(q)\Pi(i\omega_{m},q)]^{-1}V_{0}(q)\end{split}$ (S12)
We adopt $V^{RPA}(i\omega_{m}=0,q)$ as the RPA effective interaction. We can then Fourier transform to the real space:
$\begin{split}V^{RPA}(r,r^{\prime})&=\frac{1}{A}\sum_{q,g,g^{\prime}}V^{RPA}(i\omega_{m}=0,q)_{g,g^{\prime}}e^{i(q+g)r}e^{-i(q+g^{\prime})r^{\prime}}\\ &=\frac{1}{A}\sum_{q,g,g^{\prime}}V^{RPA}(i\omega_{m}=0,q)_{g,g^{\prime}}e^{iq(r-r^{\prime})}e^{igr}e^{-ig^{\prime}r^{\prime}}\end{split}$ (S13)
where it should be emphasize that, the momentum space cutoff of the above Fourier transformation should be taken relatively large, since the dual-gate screened interaction $V_{0}(q)$ in Eq. (S2) has slow $1/q$ decay at large $q$.
Fig. S5 illustrates the spatial profile of the RPA effective interaction $V^{RPA}(r,r^{\prime})$ along the NNN or NN directions. Here, $r=0$ is located at the Wannier center of A sublattice (i.e. XM region), and the dual-gate distance is fixed at experimental relevant value $\xi=10$ nm. The point $r/a_{M}=1$ for the NNN direction (Fig. S5(a)) and the point $\sqrt{3}r/a_{M}=1$ for the NN direction (Fig. S5(b)) corresponds to the Wannier center of the NNN or NN site, respectively. These results allow us to evaluate how strongly the Coulomb repulsion is screened and where the effective attraction occurs.
First, we observe that the RPA mechanism further screens the dual-gate screened Coulomb interaction, leading to even more enhanced localization of the repulsive part. This provides further justification of retaining the onsite Hubbard repulsion $H_{U}$ only in our tight-binding model, as the longer-range repulsions are significantly suppressed.
Moreover, the RPA results reveal that attractive interactions emerge predominantly in the relatively local regions (though some longer-range attractions beyond this range may also occur, but their magnitudes are strongly screened compared with the local ones). Specifically, in Fig. S5(b), a strong attraction is clearly seen near the NN region $\sqrt{3}r/a_{M}\sim 1$. While in Fig. S5(a) the attraction appears in the range $0.4\lesssim r/a_{M}\lesssim 1$ (at $r/a_{M}=1$, only weak attraction occurs for relatively large $\epsilon_{r}$), the NNN interaction can still be attractive (though weaker than the NN attraction) considering the finite spatial extent of the Wannier function. These findings justify the use of relatively local attractive terms in our tight-binding model. While both NN and NNN attraction are possible, our main text mainly focus on the NNN attraction since, as shown below (Fig. S6), it induces much stronger superconductivity, which can be theoretically understood by the sublattice polarization property of the Fermi surface.
Figure S5: The RPA effective interactions $V^{RPA}(r,r^{\prime})$ with $r$ fixed at A sublattice and $r^{\prime}$ along (a) NNN A-A (b) NN A-B direction with dual-gate distance $\xi=10$ nm and different dielectric constants $\epsilon_{r}$.
VI Derivations and Supplemental Results of Superconducting Mean-Field Analysis